The average score in the final exam of a course is 65 and the standard deviation is 10. a) Give an upper bound on the probability of a student scoring more than 95? b) Suppose the scores follow a normal distribution. Compute the probability of a student scoring more than 95 and compare it to the bound obtained in a)
Answer #1

A) less than 0.003

B) P(score > 95) = 1-p(Z < (95-65)/10)= 1- p(Z< 3) = 1-0.9987= 0.0013